OpenStudy (anonymous):

solve equation by factoring 10x-21=10/x

OpenStudy (anonymous):

x = -2/5 x = 5/2

OpenStudy (anonymous):

There are two answers.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No prob :)

OpenStudy (anonymous):

how did you factor 10xsquared-2x-10?